OpenStudy (anonymous):

(x+8)/(x^2(x^2+9)) solve using parrial fraction decompposition

OpenStudy (tkhunny):

\(\dfrac{A}{x} + \dfrac{B}{x^{2}} + \dfrac{Cx+D}{x^{2}+9}\)

OpenStudy (anonymous):

what is the error

OpenStudy (tkhunny):

?? That is the form of the partial fraction decomposition. Do you see why? You must find the coefficients.